OpenStudy (anonymous):

Explain why (–4x)^0 = 1, but –4x^0 = –4.

OpenStudy (anonymous):

X^0=1 -4*x^0=-4*1=-4

OpenStudy (anonymous):

since anything to the power of zero is 1, so in the second one only the x is raised to the zero...

OpenStudy (anonymous):

Order of operations. Powers happen before multiplication. The first expression uses parenthesis to force the multiplication to happen before exponentiation.